# TADA: Improved Diffusion Sampling with Training-free Augmented DynAmics

### v1

@@ 1 Introduction
Sampling from a diffusion model is typically discretized through a multi-step noise schedule. Denoising Diffusion Probabilistic Models sample new noise at every step resulting in a process interpretable as discretized solutions of diffusion Stochastic Differential Equations (SDEs). Meanwhile, Denoising Diffusion Implicit Models only sample initial noise which is reused at every step resulting in a process interpretable as discretized solutions of Ordinary Differential Equations (ODEs). Both ODE and SDE solvers should perform similarly since they merely represent different interpretations of the same Fokker-Planck Partial Differential Equation. In practice, however, ODE solvers tend to yield lower fidelity results compared to SDE solvers, as confirmed by FID scores in [17]. On the other hand, SDE solvers require finer steps due to the noise injection at every step, which gets amplified through discretization.
@@
    We prove the training equivalence between momentum diffusion models [9, 6] and conventional diffusion models, modulo a transformation of the input variables to the neural network.
@@
    We propose a new sampling method named Training-free Augmented DynAmics (TADA) that uses higher-dimensional input noise and yet enables direct reuse of pretrained diffusion models.
@@
    We observe that our proposed method inherits beneficial SDE properties while utilizing ODE solvers, enabling both diverse generation and accelerated sampling.
@@ 2 Preliminary
Figure 1: Here we show the distinctions between conventional diffusion models (left) and momentum diffusion models (right) during sampling. Leveraging Prop 3.1, we demonstrate that pretrained diffusion models $x_{\theta}(\cdot,\cdot)$ can be directly applied to propagate the momentum system. Moreover, the choice of numerical solver remains flexible. See Sec 3.2 for details.
@@ 2.2 Phase Space Momentum Diffusion Model
where ${\bm{A}}_{t}\in\mathbb{R}^{2\times 2}$ and ${\mathbf{b}}_{t}\in\mathbb{R}^{2}$ denote time-dependent coefficients. Specifically, in the MDM framework, the function $F_{\theta}:(\mathbb{R}^{d})^{2}\times[0,1]\rightarrow\mathbb{R}^{d}$ accepts an aggregated input of two variables of dimension $d$ and outputs a single vector of dimension $d$. The resulting $d$-dimensional output is subsequently broadcasted across the $N=2$ variables through the coefficient ${\mathbf{b}}_{t}$, a key distinction from conventional DM. The explicit formulations for ${\bm{A}}_{t}$ and ${\mathbf{b}}_{t}$ are detailed in Appendix.E.1.
@@ 2.3 Sampling with Exponential Integrators
To improve the accuracy of approximating the nonlinear integral in the second term, advanced ODE solvers are employed as approximators $\Psi(\cdot,\cdot,\cdot,\cdot)$ in eq.5. Higher-order single-step methods such as Heun’s method [12] and multi-step explicit methods like Adams–Bashforth [5] have been widely adopted in prior works [17, 23, 33, 22], often in conjunction with exponential integrators. To further enhance accuracy, implicit schemes such as the Adams–Moulton method [2] have also been introduced [34]. The current state-of-the-art solvers integrate these various techniques, and the specific combinations along with the resulting algorithms are summarized in Appendix.E.3.
@@ 3 Method
2:${\mathbf{x}}_{t_{0}}\sim\mathcal{N}(0,{\bm{\Sigma}}_{t_{0}})$triangleright draw prior sample
3:for $i=0$ to $T-1$ do
4:     compute ${\bm{\mu}}_{t_{i}},{\bm{\Sigma}}_{t_{i}},\text{and }{\mathbf{r}}_{t_{i}}:=\frac{{\bm{\Sigma}}_{t_{i}}^{-1}{\bm{\mu}}_{t_{i}}}{{\bm{\mu}}_{t_{i}}^{\mathsf{T}}{\bm{\Sigma}}_{t_{i}}^{-1}{\bm{\mu}}_{t_{i}}}$triangleright obtain the reweighting for $N$ variables
5:     $\hat{x}\leftarrow x_{\theta}\!\left(({\mathbf{r}}_{t_{i}}^{\!\top}\!\otimes I_{d}){\mathbf{x}}_{t_{i}},t_{i}\right)$triangleright data prediction with pretrained DM
6:     $F_{\theta}\leftarrow N!\dfrac{\hat{x}-\sum_{n=0}^{N-1}\frac{x_{t_{i}}^{(n)}}{n!}(1-t_{i})^{n}}{(1-t_{i})^{N}}$triangleright Compute force term, see Sec.3.2
@@
[…]
9:     ${\mathbf{x}}_{t_{i+1}}\leftarrow\Phi(t_{i+1},t_{i}){\mathbf{x}}_{t_{i}}+\Psi_{t_{i}}$triangleright state update
@@
11:return $x_{\theta}\!\bigl{(}({\mathbf{r}}_{t_{T}}^{\!\top}\!\otimes I_{d})\,{\mathbf{x}}_{t_{T}},\,t_{T}\bigr{)}$triangleright return results with data prediction
@@ 3.1 Training equivalence between Momentum Diffusion Models and Diffusion Models
Moreover, the $F_{\theta}$ in eq.4 can be recovered as a linear combination of $x_{\theta}$ and ${\mathbf{x}}_{t}$. Here, ${\bm{\mu}}_{t}^{\mathsf{T}}\mathbb{\Sigma}_{t}^{-1}{\bm{\mu}_{t}}$ is the effective SNR of $({\mathbf{r}}_{t}^{\mathsf{T}}\otimes\mathbb{I}_{d}){\mathbf{x}}_{t}$ which simply is a weighted linear combination of ${\mathbf{x}}_{t}$ by ${\mathbf{r}}_{t}$.
@@ 3.2 Momentum Diffusion Sampling Methodology
When $N=1$, this formulation simplifies precisely to vanilla flow matching. When $N=2$, it is essentially same as the deterministic case of AGM[6] but with the added training-free property that our method carries. Please see Appendix.E.4 for details.
Figure 3: Generated samples under varying prior scales are shown, all initialized with the same initial condition of dynamics in Prop.3.3:$y_{0}:=({\mathbf{r}}_{0}^{\mathsf{T}}\otimes{\bm{I}}_{d}){\mathbf{x}}_{0}\equiv\epsilon$, using an identical time discretization over the SNR and the same pretrained model with 15 NFEs. It can be observed that the diversity of the generated results increases proportionally with the standard deviation of the final variable $x_{0}^{(N-1)}$, which is scaled by a factor $k$: ${\bm{\Sigma}}_{0}=\text{diag}(1,1,...,k)$.
To compute the reweighting term ${\mathbf{r}}_{t}$, the mean and covariance matrix at time $t$ are required. These quantities can be obtained by analytically solving the coupled dynamics of the mean ${\bm{\mu}_{t}}$ and covariance ${\bm{\Sigma}_{t}}$, a standard approach [26] in both conventional DM and MDM. Due to space constraints, the explicit analytical expressions are presented in Appendix E.2.
@@ 3.3 Analysis of Sampling Dynamics
For the coefficient of $w_{t}^{i}$, $\alpha_{t}$ and $\beta_{t}$, please refer to Appendix.B.
@@
There are two scenarios in which Prop 3.3 degenerates into a mere different time discretization of the conventional diffusion model, irrespective of the value of $N$. The first occurs when $N=1$, and the second arises when ${\mathbf{A}}_{t}$ is a diagonal matrix, implying that each variable evolves independently. Further details and proof are provided in Appendix.E.5.
@@ 4 Experiments
Figure 4: Qualitative comparison with UniPC, varying the NFEs, using the same initial condition and the same pretrained EDM2 model. More qualitative comparision can be found in Appendix.F.
@@
Figure 5: left: Comparison with baselines on ImageNet-64 using EDM pretrained model. Right: Performance under varying numbers of variables $N$ while keeping the SNR-based time discretization same to the $N=2$ setting;$N=2$ is the default configuration reported throughout this paper.
@@ 4.1 EDM and EDM2 Experiment
We evaluate performance using Fréchet Inception Distance score(FID;Heusel et al. [13]) for both ImageNet-64 and ImageNet-512, and additionally use Fréchet Distance-DINOv2(FD-DINOv2;Stein et al. [29]) for ImageNet-512. Our results show that TADA consistently outperforms the baselines across all tested numbers of function evaluations (NFEs) in fig.5 and fig.6. We also examine the case with $N>2$, which introduces additional pseudo noise as described in Prop.3.3. In all setups, we use the same discretization over SNR and vary only the standard deviation of $x_{0}^{(N-1)}$. This configuration results in improved performance on ImageNet-64, as shown in fig 5. However, no consistent improvement is observed on ImageNet-512, which may be attributed to the high capacity of the neural network and limited exploration of time discretization, thereby diminishing the impact of the additional noise perturbation. For consistency, we therefore report results only for the $N=2$ in fig.6 and the rest of experiments.
Fig 4 presents qualitative results under varying NFE budgets. Unlike conventional ODE solvers, which tend to produce increasingly blurred outputs as NFEs decrease, TADA continues to generate plausible images with preserved details. Nevertheless, these images may still deviate from the data distribution $p_{\rm{data}}$, as indicated by the larger FID metric. See Appendix.F for more examples.
@@ 5 Conclusions and Limitations
In this appendix, we follow the notation introduced before. However, in order to reduce the complexity of the derivation, we simplify all the coefficient with Kronecker product. For example ${\mathbf{A}}_{t}:={\mathbf{A}}_{t}\otimes{\mathbf{I}}_{d}$, ${\bm{\mu}_{t}}:={\bm{\mu}_{t}}\otimes{\mathbf{I}}_{d}$.
Appendix A Proof of Proposition.3.1
@@
We first analyze the objective function of momentum diffusion model for $N=2$ case, and it can be generalize to larger $N$.
@@
Following the same spirit, one can derive the case for $N$ variable. See Appendix.I for details.
@@
Appendix B Proof of Proposition.3.3
@@
Consider the linear time-varying system:
@@
Since (26) is linear and deterministic (apart from the random initial condition), the state remains Gaussian. Its mean and covariance evolve as follows [26].
@@
Using (30) and $\dot{{\bm{\mu}_{t}}}=\hat{A}_{t}{\bm{\mu}_{t}}+\hat{{\mathbf{b}}}_{t}$, one obtains (after cancellation) the result:
@@
The frist term is essentially one kind of linear combination of ${\mathbf{x}}_{t}$, and Recally that $y_{t}={\mathbf{r}}^{\mathsf{T}}{\mathbf{x}}_{t}:=\frac{{\bm{\mu}}_{t}^{\mathsf{T}}{\bm{\Sigma}}_{t}^{-1}}{{\bm{\mu}}_{t}^{\mathsf{T}}{\bm{\Sigma}}_{t}^{-1}{\bm{\mu}_{t}}}$ which is another linear combination of ${\mathbf{x}}_{t}$. Assume that ${\mathbf{x}}_{t}\sim\mathcal{N}({\bm{\mu}_{t}}x_{1},{\bm{\Sigma}}_{t})$, thus, one can derive the relationship between the frist term and $y_{t}$. Thus, according to Lemma.H.1
@@
ϵsubscriptbold-italic-${\bm{\epsilon}}_{\perp}$ $\sim\mathcal{N}\!\bigl{(}\mathbf{0},\ I_{d}-{\mathbf{L}}_{t}^{\mathsf{T}}{\mathbf{r}}_{t}{\mathbf{r}}_{t}^{\mathsf{T}}{\mathbf{L}}_{t}/{\mathbf{r}}_{t}^{\mathsf{T}}{\bm{\Sigma}}_{t}{\mathbf{r}}_{t}\bigr{)}.$ (41)
@@
Appendix C Experiment Details
@@ C.1 EDM and EDM2
[…]
@@ C.2 Stable Diffusion 3
Appendix D Additional Plots
D.1 General $N$ variable dynamics

### v2

@@ 1 Introduction
Sampling from a diffusion model is typically discretized through a multi-step noise schedule. Denoising Diffusion Probabilistic Models sample new noise at every step resulting in a process interpretable as discretized solutions of diffusion Stochastic Differential Equations (SDEs). Meanwhile, Denoising Diffusion Implicit Models only sample initial noise which is reused at every step resulting in a process interpretable as discretized solutions of Ordinary Differential Equations (ODEs). Both ODE and SDE solvers should perform similarly since they merely represent different interpretations of the same Fokker-Planck Partial Differential Equation. In practice, however, ODE solvers often yield lower‐fidelity results than SDE solvers as evidenced by the FID scores in  [17] when sufficient function evaluations are performed and model capacity is a limiting factor. On the other hand, SDE solvers require finer steps due to the noise injection at every step, which gets amplified through discretization.
@@
    We prove the training equivalence between momentum diffusion models  [9, 6] and conventional diffusion models, modulo a transformation of the input variables to the neural network. Consequently, the equivalence shows that simple noise or noisy data augmentation together with diffusion loss, including but not limited to momentum diffusion, offers no training benefit. Readers may extend this reasoning to their own settings and verify the algorithms relevant to them.
@@
    We then shift our attention to the sampling phase of the momentum system. We propose a new sampling method named Training-free Augmented DynAmics (TADA) that uses higher-dimensional input noise and yet enables direct reuse of pretrained diffusion models.
@@
    Although no benefits are observed in the training phase in theory, we find that our proposed method, rooted in the momentum system, inherits advantageous SDE properties while employing ODE solvers, enabling both diverse generation and accelerated sampling through momentum dynamics.
@@ 2 Preliminary
Figure 1: Here we show the distinctions between conventional diffusion models (left) and momentum diffusion models (right) during sampling. Leveraging Prop 3.1, we demonstrate that pretrained diffusion models with $x_{0}$-prediction $x_{\theta}(\cdot,\cdot)$ can be directly applied to propagate the momentum system with multiple varible $\{{\mathbf{x}}_{t_{i}}^{(}n)\}_{0}^{N-1}$. Moreover, the choice of numerical solver remains flexible. See Sec 3.2 for details.
@@ 2.2 Phase Space Momentum Diffusion Model
where ${\bm{A}}_{t}\in\mathbb{R}^{2\times 2}$ and ${\mathbf{b}}_{t}\in\mathbb{R}^{2}$ denote time-dependent coefficients. Specifically, in the MDM framework, the function $F_{\theta}:(\mathbb{R}^{d})^{2}\times[0,1]\rightarrow\mathbb{R}^{d}$ accepts an aggregated input of two variables of dimension $d$ and outputs a single vector of dimension $d$. The resulting $d$-dimensional output is subsequently broadcasted across the $N=2$ variables through the coefficient ${\mathbf{b}}_{t}$, a key distinction from conventional DM. The explicit formulations for ${\bm{A}}_{t}$ and ${\mathbf{b}}_{t}$ are detailed in Appendix.F.1.
@@ 2.3 Sampling with Exponential Integrators
To improve the accuracy of approximating the nonlinear integral in the second term, advanced ODE solvers are employed as approximators $\Psi(\cdot,\cdot,\cdot,\cdot)$ in eq.5. Higher-order single-step methods such as Heun’s method [12] and multi-step explicit methods like Adams–Bashforth [5] have been widely adopted in prior works [17, 23, 33, 22], often in conjunction with exponential integrators. To further enhance accuracy, implicit schemes such as the Adams–Moulton method [2] have also been introduced [34]. The current state-of-the-art solvers integrate these various techniques, and the specific combinations along with the resulting algorithms are summarized in Appendix.F.3.
@@ 3 Method
3:for $i=0$ to T1T-1 do
4:  compute ${\bm{\mu}}_{t_{i}},{\bm{\Sigma}}_{t_{i}},\text{and }{\mathbf{r}}_{t_{i}}:=\frac{{\bm{\Sigma}}_{t_{i}}^{-1}{\bm{\mu}}_{t_{i}}}{{\bm{\mu}}_{t_{i}}^{\mathsf{T}}{\bm{\Sigma}}_{t_{i}}^{-1}{\bm{\mu}}_{t_{i}}}$ $\triangleright$ obtain the reweighting for NN variables
5:  $\hat{x}\leftarrow x_{\theta}\!\left(({\mathbf{r}}_{t_{i}}^{\!\top}\!\otimes I_{d}){\mathbf{x}}_{t_{i}},t_{i}\right)$ $\triangleright$ data prediction with pretrained DM
6:  $F_{\theta}\leftarrow N!\dfrac{\hat{x}-\sum_{n=0}^{N-1}\frac{x_{t_{i}}^{(n)}}{n!}(1-t_{i})^{n}}{(1-t_{i})^{N}}$ $\triangleright$ Compute force term, see Sec. 3.2
7:  $\Psi_{t_{i}}\approx\displaystyle\int_{t_{i}}^{t_{i+1}}\!\!\Phi(t_{i+1},\tau)\,{\mathbf{b}}_{\tau}F({\mathbf{x}}_{\tau},\tau)\,d\tau$ $\triangleright$ Approx. nonliear part using existing solver (with $Q$ if solver is multistep)
8:  if Solver is multistep then $Q\xleftarrow{\text{cache}}\hat{x} \triangleright$ store history
9:  ${\mathbf{x}}_{t_{i+1}}\leftarrow\Phi(t_{i+1},t_{i}){\mathbf{x}}_{t_{i}}+\Psi_{t_{i}}$ $\triangleright$ state update
@@
11:return $x_{\theta}\!\bigl(({\mathbf{r}}_{t_{T}}^{\!\top}\!\otimes I_{d})\,{\mathbf{x}}_{t_{T}},\,t_{T}\bigr)$ $\triangleright$ return results with data prediction
@@ 3.1 Training equivalence between Momentum Diffusion Models and Diffusion Models
Moreover, the $F_{\theta}$ in eq. 4 can be recovered as a linear combination of $x_{\theta}$ and ${\mathbf{x}}_{t}$ (see section.3.2 for details). Here, ${\bm{\mu}}_{t}^{\mathsf{T}}\mathbb{\Sigma}_{t}^{-1}{\bm{\mu}_{t}}$ is the effective SNR of $({\mathbf{r}}_{t}^{\mathsf{T}}\otimes\mathbb{I}_{d}){\mathbf{x}}_{t}$ which simply is a weighted linear combination of ${\mathbf{x}}_{t}$ by ${\mathbf{r}}_{t}$.
@@ 3.2 Momentum Diffusion Sampling Methodology
When $N=1$, this formulation simplifies precisely to vanilla flow matching. When $N=2$, it is essentially same as the deterministic case of AGM [6] but with the added training-free property that our method carries. Please see Appendix. F.4 for details.
[…]
To compute the reweighting term ${\mathbf{r}}_{t}$, the mean and covariance matrix at time $t$ are required. These quantities can be obtained by analytically solving the coupled dynamics of the mean $\bm{\mu}_{t}$} and covariance $\bm{\Sigma}_{t}$}, a standard approach  [26] in both conventional DM and MDM. Due to space constraints, the explicit analytical expressions are presented in Appendix F.2.
@@ 3.3 Analysis of Sampling Dynamics
For the coefficient of $w_{t}^{i}$, $\alpha_{t}$ and $\beta_{t}$, please refer to Appendix. C. $\epsilon_{t}^{(i)}$ denotes the estimated Gaussian noise for each variable $x_{t}^{i}$, induced by $x_{\theta}$. This is analogous to the standard diffusion model but extended to the multi-variable case. Please see Appendix.C for more detail.
@@
There are two scenarios in which Prop 3.3 degenerates into a mere different time discretization of the conventional diffusion model, irrespective of the value of $N$. The first occurs when $N=1$, and the second arises when ${\mathbf{A}}_{t}$ is a diagonal matrix, implying that each variable evolves independently. Further details and proof are provided in Appendix. F.5.
@@ 4 Experiments
Figure 4: Qualitative comparison with UniPC, varying the NFEs, using the same initial condition and the same pretrained EDM2 model. More qualitative comparision can be found in Appendix. G.
@@
Figure 5: left: Comparison with baselines on ImageNet-64 using EDM pretrained model. Right: Performance under varying numbers of variables NN while keeping the SNR-based time discretization same to the $N=2$ setting;$N=2$ is the default configuration reported throughout this paper.
@@ 4.1 EDM and EDM2 Experiment
We evaluate performance using Fréchet Inception Distance score(FID [13]) for both ImageNet-64 and ImageNet-512, and additionally use Fréchet Distance-DINOv2 (FD-DINOv2 [29]) for ImageNet-512. Our results show that TADA consistently outperforms the baselines across all tested numbers of function evaluations (NFEs) in fig. 5 and fig. 6. We also examine the case with $N>2$, which introduces additional pseudo noise as described in Prop. 3.3. In all setups, we use the same discretization over SNR and vary only the standard deviation of $x_{0}^{(N-1)}$. This configuration results in improved performance on ImageNet-64, as shown in fig 5. However, no consistent improvement is observed on ImageNet-512, which may be attributed to the high capacity of the neural network and limited exploration of time discretization, thereby diminishing the impact of the additional noise perturbation. For consistency, we therefore report results only for the $N=2$ in fig. 6 and the rest of experiments.
Fig 4 presents qualitative results under varying NFE budgets. Unlike conventional ODE solvers, which tend to produce increasingly blurred outputs as NFEs decrease, TADA continues to generate plausible images with preserved details. Nevertheless, these images may still deviate from the data distribution $p_{\rm{data}}$, as indicated by the larger FID metric. See Appendix. G for more examples.
@@ 5 Conclusions and Limitations
Appendix A Notations
In this appendix, we follow the notation introduced before. However, in order to reduce the complexity of the derivation, we simplify all the coefficient with Kronecker product. All the variable which sololy depends on the time, will be broadcast by the kroneker product, For example ${\mathbf{A}}_{t}:={\mathbf{A}}_{t}\otimes{\mathbf{I}}_{d}$, ${\bm{\mu}_{t}}:={\bm{\mu}_{t}}\otimes{\mathbf{I}}_{d}$, ${\mathbf{L}}_{t}^{v}v={\mathbf{L}}_{t}^{v}v\otimes{\mathbf{I}}_{d}$ etc.
Appendix B Proof of Proposition. 3.1
@@ Appendix A Notations
We first analyze the objective function of momentum diffusion model for $N=2$ case, and it can be generalize to larger $N$. For clarity in dimensional alignment and derivational correctness, we refer the reader to the simplified notation in Section A, which will also be adopted throughout this section.
@@
Following the same spirit, one can derive the case for $N$ variable. See Appendix. J for details.
@@
Appendix C Proof of Proposition. 3.3
For clarity in dimensional alignment and derivational correctness, we refer the reader to the simplified notation in Section A, which will also be adopted throughout this section.
@@
Consider the linear time‐varying system:
@@
Since ( 26) is linear and deterministic (apart from the random initial condition), the state remains Gaussian. Its mean and covariance evolve as follows  [26].
@@
Using ( 30) and $\dot{{\bm{\mu}_{t}}}=\hat{A}_{t}{\bm{\mu}_{t}}+\hat{{\mathbf{b}}}_{t}$, one obtains (after cancellation) the result:
@@
The frist term is essentially one kind of linear combination of ${\mathbf{x}}_{t}$, and Recally that $y_{t}={\mathbf{r}}^{\mathsf{T}}{\mathbf{x}}_{t}:=\frac{{\bm{\mu}}_{t}^{\mathsf{T}}{\bm{\Sigma}}_{t}^{-1}}{{\bm{\mu}}_{t}^{\mathsf{T}}{\bm{\Sigma}}_{t}^{-1}{\bm{\mu}_{t}}}$ which is another linear combination of ${\mathbf{x}}_{t}$. Assume that ${\mathbf{x}}_{t}\sim\mathcal{N}({\bm{\mu}_{t}}x_{1},{\bm{\Sigma}}_{t})$, thus, one can derive the relationship between the frist term and $y_{t}$. Thus, according to Lemma. I.1
@@
${\bm{\epsilon}}_{\perp}$ $\sim\mathcal{N}\!\bigl(\mathbf{0},\ I_{d}-{\mathbf{L}}_{t}^{\mathsf{T}}{\mathbf{r}}_{t}{\mathbf{r}}_{t}^{\mathsf{T}}{\mathbf{L}}_{t}/{\mathbf{r}}_{t}^{\mathsf{T}}{\bm{\Sigma}}_{t}{\mathbf{r}}_{t}\bigr).$ (41)
@@
Appendix D Experiment Details
@@ D.1 EDM and EDM2
For DPM-Solver++, we did abalation search over $\text{order}\in[1,2,3]$, $\text{discretization}\in[\text{logSNR},\text{ time uniform},,\text{edm},\text{time quadratic}]$.
For UniPC, we did abalation search over $\text{order}\in[1,2,3]$, $\text{discretization}\in[\text{logSNR},\text{ time uniform},,\text{edm},\text{time quadratic}]$$\text{variant}\in[\text{bh1},\text{bh2}]$.
@@ D.2 Stable Diffusion 3
Appendix E Additional Plots
E.1 General $N$ variable dynamics